OpenStudy (anonymous):

Which statement correctly compares 11/20 and 4/7 A.11/20=4/7 B.11/20>4/7 c 11/20<4/7

OpenStudy (anonymous):

c 11/20<4/7

OpenStudy (anonymous):

thanks

OpenStudy (helder_edwin):

\[\large \frac{a}{b}=\frac{c}{d}\quad\Leftrightarrow\quad ad=bc \] \[\large \frac{a}{b}<\frac{c}{d}\quad\Leftrightarrow\quad ad<bc \]

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

lol